OpenStudy (anonymous):

lim x→0 (x^2− 2 sin x)/x What do? I'm not a big fan of limits, so all help is appreciated.

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Use L'Hopital's Rule

OpenStudy (anonymous):

\[ \lim_{x\rightarrow 0} \frac{x^2-2\sin(x)}{x} = \lim_{x\rightarrow 0} x - 2\cdot \lim_{x\rightarrow 0} \frac{\sin(x)}{x} \]

OpenStudy (anonymous):

Now the first of the two limits on the right evaluates to 0, clearly.

OpenStudy (anonymous):

And for the second, you can use L'Hopitals Rule or the Taylor expansion of sine, as you wish.

OpenStudy (anonymous):

By L'Hopitals Rule you get \[ \lim_{x\rightarrow 0} \frac{\sin(x)}{x} = \lim_{x\rightarrow 0} \frac{(\sin(x))^\prime}{(x)^\prime} = \lim_{x\rightarrow 0} \frac{\cos(x)}{1} = 1 \] where the ' denotes derivative with respect to x and the last equality holds because \[\cos(0)=1 \]

OpenStudy (anonymous):

Putting it all together your limit evaluates to -2.

OpenStudy (anonymous):

Thank you that actually clears a lot of this limit thing up.

OpenStudy (anonymous):

Wonderful Manifold

OpenStudy (zarkon):

\[\lim_{x\rightarrow 0} \frac{\sin(x)}{x} = 1\] is one limit that should be memorized. To see why it is true just look at any undergradute text on calculus.

myininaya (myininaya):

\[\lim_{x \rightarrow 0}(\frac{x^2}{x}-2 \cdot \frac{\sin(x)}{x})=\lim_{x \rightarrow 0}x-2 \lim_{x \rightarrow 0}\frac{\sin(x)}{x}\] \[0-2(1)=-2\]

myininaya (myininaya):

It is true by squeeze thm or sandwich thm

myininaya (myininaya):

whatever you want to call it

OpenStudy (zarkon):

I usually go by squeeze theorem myself :)

myininaya (myininaya):

i like sandwiches though

OpenStudy (zarkon):

sandwiches are good...but i don't want to be thinking about food when I'm doing math

myininaya (myininaya):

maybe that is wise